# Incentivizing News Consumption on Social Media Platforms Using Large Language Models and Realistic Bot Accounts

### v2

@@ Abstract
This manuscript was compiled on March 21, 2024
@@ 1 Introduction
We argue that the problem is less that people consume bad political information, but that most people do not consume any at all. News and politics constitute a small fraction of people’s information diets on social media. News makes up only 1.4% of Facebook’s News Feed (13, 14), the majority of Twitter users do not follow any politicians, journalists, or news organizations (6), and only about 1 in 300 outbound clicks from social media are to substantive news (15). [^1] This is the case online more broadly: only between 2% (16) and 7-9% (17) of all URLs visited by large international samples are news, and - across mobile and desktop - news comprises only 4% of total consumption (18). Although news consumption is greater on television than online, it is still overshadowed by entertainment and other content categories (19, 20)).
This under-consumption of news—and the consequently low levels of political knowledge among the electorate—have important implications (21). Low-information voters are likely to withdraw from politics, vulnerable to making irrational vote choices (22, 23, 24), and easily swung by irrelevant stimuli, emotional appeals, and populist rhetoric in the political environment (25). If sizable, these voters can swing elections (23, 26). In turn, exposure to verified and ideologically balanced quality news creates an informed and efficacious public and leads to more stable political attitudes, lower susceptibility to misinformation, greater acceptance of democratic norms, and voting in accordance with one’s personal and group interests (21, 27, 28). Some note that news exposure is the key predictor of political knowledge, seen as “a demonstrably critical foundation for good citizenship” (27).
@@
We find that encouragement to follow news through our tailored NLP-based responses had some promising yet limited causal effects; it encouraged the users to follow more news outlets and additionally encouraged those who received comments from the female bot to like more news content on social media. The treatment had no effects on the other outcomes analyzed, i.e. (re)tweeting news content and tweeting about or liking posts about politics. The increases in the liking of news media content were confined to those with high initial levels of political interest, as indicated by previous tweeting about politics, suggesting reinforcement of pre-existing engagement among those already engaged (31, 32). Also, the effects on news media content liking were especially pronounced for the users who were tweeting about sports, with the effects among those who tweeted about entertainment or lifestyle being statistically insignificant due to the decreased sample sizes in these two topic categories.
@@
This project advances past work in several key ways. First, we address the core problem of low news use and engagement (33, 34, 35). The overwhelming majority of social media users go online for entertainment, not news or politics (13, 36, 37, 38, 39, 40). Because information exposure on platforms is primarily driven by recommender algorithms that make automated decisions on what content to display based on the user’s past behavior and inferred interests (41), [^3] Recommendation algorithms drive between 75-95% of information consumption on platforms (41)., those users are mostly recommended contents about sports, movies, or celebrities. These personalized recommendations ultimately create closed loops of entertainment consumption and narrow information repertoires (42, 36). Our intervention encourages users’ to follow news organizations. This puts public affairs information in the users’ inventory, thus increasing the likelihood that users have such information readily available in their feed (43). In addition, following news accounts signals some political interest to the algorithm, thus generating subsequent recommendations to news and political content (36).
@@
Second, we propose that social media users can be encouraged to consume news and public affairs through their non-political interests. We engage users interested in sports, entertainment, and lifestyle by connecting news with these topics and directing users to news outlets that offer both hard news and softer news about the topics of interest to the users (e.g., coverage of sports, movies, or wellness or style). The work on soft news or ”infotainment” indeed suggests that programs that discuss cooking or celebrities, but also mention current affairs, attract viewers whose primary motivation is not politics, but who nevertheless learn about current affairs and become more politically active (44, 45, 46, 47, 48, 49). Accordingly, by starting from citizen’s interests in sports, entertainment, or lifestyle topics such as yoga or keto diet, and by directing them to topically relevant content in primarily hard news outlets (e.g., the lifestyle or sports sections of ABC News), we can enhance users’ interest and sustainably increase their exposure to quality news. Social media platforms have the potential to act as an intermediary to news media organizations (50, 51) and so encouraging users to follow news accounts and to visit news sites through links embedded in posts may serve as a gateway to hard news consumption (52).
@@
Third, differential engagement may occur based on who is sharing news with the users and also who the users are. In general, news and politics are seen as male-dominated spaces (53, 54) and, notably, female media figures are found to receive more toxic, abusive, and hostile responses compared to their male counterparts (55, 56, 57). This is especially true in political contexts (58) and this disparity may be most pronounced on social media (59). It is likely then that users may be less inclined to open links from female sources or to see them as a news source worth following or interacting with. In addition, there may be heterogeneous treatment effects by users’ political interest. On the one hand, those politically disinterested individuals may gain more in terms of knowledge, engagement, or subsequent news seeking from inadvertently encountering politics (46, 52), such as seeing links to news websites on social media platforms. On the other hand, inadvertent exposure to news and politics may reinforce existing gaps in prior political interest (60, 31), such that those who are already interested in news and politics would follow more news accounts and engage most with public affairs online. In short, individuals may be responsive to social media nudges but the source (gender) of these prompts as well as prior users’ posting about current affairs on the platform may be important factors influencing these results.
@@ 2 Data and Measurement
We also measure to what extent the treated users interacted with our bots, by checking whether they replied to it. Lastly, we evaluate the sentiment of the replies using a RoBERTa-base model trained on 124M tweets from January 2018 to December 2021, and fine-tuned for sentiment analysis with the TweetEval benchmark (65, 66). We collected a total of 241 (99 Male and 142 Female) responses and examined their sentiment. See SM Incentivizing News Consumption on Social Media Platforms Using Large Language Models and Realistic Bot AccountsE for details.
@@ 3.2 Treatment Effects
As our intervention was dependent on users actually tweeting about our keywords during the experimental time period, not all users in the treatment groups received the treatment from one of our bots. To account for this, we estimate two distinct treatment effects: “Intention to treat (ITT)” (i.e., all users from our original randomized treatment groups) and “Treated” (i.e., users in the male and female treatment groups who actually received one or more responses from our bots). Because certain users may have been more likely to tweet matching keywords and view the bot responses, and so simply dropping untreated units would result in imbalance between the control group and the refined treatment groups, we use an entropy balancing approach (67) to reweigh our “treated” treatment groups relative to the control group. The entropy balancing is based on four account level metrics that capture the overall size and activity levels of an account and that are also correlated with the likelihood of receiving and seeing a treatment (total likes, total tweets, total followers, and total followed accounts). After balancing, the estimates are computed using G-computation (68, 69) with robust standard errors. For comparability, estimates are standardized with effects interpreted in standard deviation changes.
@@ 4 Discussion
The American public is largely disinterested in politics (70) and so the aggregate consumption of news and political information is limited offline (33, 34, 35, 71), online (72) and on social media platforms (13, 73, 36, 43). And yet, because news consumption has many beneficial effects—ranging from increased political knowledge and participation, to more stable political attitudes, greater political tolerance, and higher support for democratic norms (21, 27, 74)—the question as to how to incentivize social media users to consume more quality news is important. Because most citizens go online for entertainment (40), this project proposed to link users’ online habits and interests with contextual nudges that could encourage them to consume more news and public affairs information on Twitter.
@@
Second, the bots presented as females led to more consistent and stronger increases in news following and the liking of news media content. Despite the fact that politics is still seen as a male domain (53, 54) and despite the mounting evidence that women are more likely than men to experience harassment or hate speech on social media (55, 56, 57), our sample was slightly more responsive to the intervention when it came from a female. Because Twitter users are predominantly male (61.2% vs 38.8%) (76) and because our sample was likely to have even more male users (given that sports was the dominant category), it is possible that men are more open to female bots nudging them and could feel potentially threatened by male bots telling them to follow news. Again, because these effects were small and not very robust, and because this evidence goes against the prevalent finding that politics and the online sphere are spaces where female opinion is disregarded, we encourage more work exploring these potentially differential reactions to pro-social online interventions coming from females versus males.
Third, the small detected effects were largely confined to the group of users who were already interested in politics. It is those who had previously tweeted about politics who showed the most substantial increases in the liking of news media contents as a result of our intervention (among those who were treated). In contrast, the bots failed to encourage users who were not interested in politics, as determined by their previous posting of political content on Twitter, to follow more news accounts, (re)tweet politics, or like news and political content.This finding speaks to the reinforcement hypothesis (31, 32) and suggests that people who are already politically interested are the ones who become yet more politically interested as a result of various interventions. Although incidental exposure to news and public affairs on social media platforms—which the comments from our bots effectively created—could serve an equalizing function and pull the previously disinterested citizens back into news and politics (46)—our study adds to the more pessimistic evidence that individuals with high political interest are becoming information richer and more participatory, whereas the ones with low political interest remain politically disengaged (31). We also note that whatever significant effects mostly emerged among the users who were tweeting about sports, potentially due to the fact that sports users were the majority of our sample. We caution against putting too much leverage on these findings because the sample sizes became substantially smaller when sub-divided into the three topic categories.
@@
In addition, it is not certain whether similar effects would emerge on a different platform, in a different time, or in a different sociopolitical context. Twitter is known to be an important channel for political information (77, 78), a key platform for politicians, journalists, and pundits (79), and one where many users express their political opinions and engage in political activities (80, 81). As such, engaging users with news and politics may have been more “natural” on and better integrated with Twitter than with Facebook or Instagram, where many users have more closely knit networks of friends and family. We encourage scholars to replicate our results on other platforms. Given the changes to the Twitter API after Elon Musk acquired Twitter, field experiments such as ours and others (82, 83, 30) might be no longer possible.
Lastly, our core focus was on encouraging the users to follow news accounts and engage with news and political information on social media, and so we cannot ascertain whether the treatment or the slightly increased news following and liking of news content had any effects on users’ political attitudes. A growing body of evidence suggest that although various (algorithmic) interventions can powerfully alter users’ on-platform exposures and behaviors, this has no corresponding effects on affective polarization, misperceptions, policy positions, among other attitudes (43, 84, 85, 86, 75). As aforementioned, however, our treatment could have triggered or enhanced political interest among some users, served as a gateway to hard news, or made users feel politically efficacious, outcomes that we did not measure and that are often overlooked in similar work.
@@
The API call we used to get the followed accounts was API.get_friends_ids(). The call for likes was API.get_favorites() and the call for (re-)tweets was API.user_timeline(). The final counts collected were as follows: followed accounts (pre N = 6,536,692, post N = 17,286,211), (re-)tweets (pre N = 2,285,401, post N = 2,201,009), and likes (pre N = 2,927,951, post N 2,846,354). More details can be found in SM Incentivizing News Consumption on Social Media Platforms Using Large Language Models and Realistic Bot Accounts
@@
The authors gratefully acknowledge the support of the European Research Council, “Europeans exposed to dissimilar views in the media: investigating backfire effects,” Proposal EXPO- 756301 (ERC Starting Grant, Magdalena Wojcieszak – PI). The authors are also grateful to Dr Zubair Shafiq and Muhammad Haroon for their research assistance. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the European Research Council.

### v3

@@ Abstract
This manuscript was compiled on May 2, 2024
@@ 1 Introduction
We argue that the problem is less that people consume bad political information, but that most people do not consume any at all. News and politics constitute a small fraction of people’s information diets on social media. News makes up only 1.4% of Facebook’s News Feed (13, 14), the majority of Twitter users do not follow any politicians, journalists, or news organizations (6), and only about 1 in 300 outbound clicks from social media are to substantive news (15). [^1] This is the case online more broadly: only between 2% (16) and 7-9% (17) of all URLs visited by large international samples are news, and - across mobile and desktop - news comprises only 4% of total consumption (18). Although news consumption is greater on television than online, it is still overshadowed by entertainment and other content categories (18)).
This under-consumption of news—and the consequently low levels of political knowledge among the electorate—have important implications (19). Low-information voters are likely to withdraw from politics, vulnerable to making irrational vote choices (20, 21, 22), and easily swung by irrelevant stimuli, emotional appeals, and populist rhetoric in the political environment (23). If sizable, these voters can swing elections (21, 24). In turn, exposure to verified and ideologically balanced quality news creates an informed and efficacious public and leads to more stable political attitudes, lower susceptibility to misinformation, greater acceptance of democratic norms, and voting in accordance with one’s personal and group interests (19, 25, 26). Some note that news exposure is the key predictor of political knowledge, seen as “a demonstrably critical foundation for good citizenship” (25).
@@
We find that encouragement to follow news through our tailored NLP-based responses had some promising yet limited causal effects; it encouraged the users to follow more news outlets and additionally encouraged those who received comments from the female bot to like more news content on social media. The treatment had no effects on the other outcomes analyzed, i.e. (re)tweeting news content and tweeting about or liking posts about politics. The increases in the liking of news media content were confined to those with high initial levels of political interest, as indicated by previous tweeting about politics, suggesting reinforcement of pre-existing engagement among those already engaged (29, 30). Also, the effects on news media content liking were especially pronounced for the users who were tweeting about sports, with the effects among those who tweeted about entertainment or lifestyle being statistically insignificant due to the decreased sample sizes in these two topic categories.
@@
This project advances past work in several key ways. First, we address the core problem of low news use and engagement (31, 32, 33). The overwhelming majority of social media users go online for entertainment, not news or politics (13, 34, 35, 36, 37, 38). Because information exposure on platforms is primarily driven by recommender algorithms that make automated decisions on what content to display based on the user’s past behavior and inferred interests (39), [^3] Recommendation algorithms drive between 75-95% of information consumption on platforms (39)., those users are mostly recommended contents about sports, movies, or celebrities. These personalized recommendations ultimately create closed loops of entertainment consumption and narrow information repertoires (40, 34). Our intervention encourages users’ to follow news organizations. This puts public affairs information in the users’ inventory, thus increasing the likelihood that users have such information readily available in their feed (41). In addition, following news accounts signals some political interest to the algorithm, thus generating subsequent recommendations to news and political content (34).
@@
Second, we propose that social media users can be encouraged to consume news and public affairs through their non-political interests. We engage users interested in sports, entertainment, and lifestyle by connecting news with these topics and directing users to news outlets that offer both hard news and softer news about the topics of interest to the users (e.g., coverage of sports, movies, or wellness or style). The work on soft news or ”infotainment” indeed suggests that programs that discuss cooking or celebrities, but also mention current affairs, attract viewers whose primary motivation is not politics, but who nevertheless learn about current affairs and become more politically active (42, 43, 44, 45, 46, 47). Accordingly, by starting from citizen’s interests in sports, entertainment, or lifestyle topics such as yoga or keto diet, and by directing them to topically relevant content in primarily hard news outlets (e.g., the lifestyle or sports sections of ABC News), we can enhance users’ interest and sustainably increase their exposure to quality news. Social media platforms have the potential to act as an intermediary to news media organizations (48, 49) and so encouraging users to follow news accounts and to visit news sites through links embedded in posts may serve as a gateway to hard news consumption (50).
@@
Third, differential engagement may occur based on who is sharing news with the users and also who the users are. In general, news and politics are seen as male-dominated spaces (51, 52) and, notably, female media figures are found to receive more toxic, abusive, and hostile responses compared to their male counterparts (53, 54, 55). This is especially true in political contexts (56) and this disparity may be most pronounced on social media (57). It is likely then that users may be less inclined to open links from female sources or to see them as a news source worth following or interacting with. In addition, there may be heterogeneous treatment effects by users’ political interest. On the one hand, those politically disinterested individuals may gain more in terms of knowledge, engagement, or subsequent news seeking from inadvertently encountering politics (44, 50), such as seeing links to news websites on social media platforms. On the other hand, inadvertent exposure to news and politics may reinforce existing gaps in prior political interest (58, 29), such that those who are already interested in news and politics would follow more news accounts and engage most with public affairs online. In short, individuals may be responsive to social media nudges but the source (gender) of these prompts as well as prior users’ posting about current affairs on the platform may be important factors influencing these results.
@@ 2 Data and Measurement
We also measure to what extent the treated users interacted with our bots, by checking whether they replied to it. Lastly, we evaluate the sentiment of the replies using a RoBERTa-base model trained on 124M tweets from January 2018 to December 2021, and fine-tuned for sentiment analysis with the TweetEval benchmark (63, 64). We collected a total of 241 (99 Male and 142 Female) responses and examined their sentiment. See SM Incentivizing News Consumption on Social Media Platforms Using Large Language Models and Realistic Bot AccountsE for details.
@@ 3.2 Treatment Effects
As our intervention was dependent on users actually tweeting about our keywords during the experimental time period, not all users in the treatment groups received the treatment from one of our bots. To account for this, we estimate two distinct treatment effects: “Intention to treat (ITT)” (i.e., all users from our original randomized treatment groups) and “Treated” (i.e., users in the male and female treatment groups who actually received one or more responses from our bots). Because certain users may have been more likely to tweet matching keywords and view the bot responses, and so simply dropping untreated units would result in imbalance between the control group and the refined treatment groups, we use an entropy balancing approach (65) to reweigh our “treated” treatment groups relative to the control group. The entropy balancing is based on four account level metrics that capture the overall size and activity levels of an account and that are also correlated with the likelihood of receiving and seeing a treatment (total likes, total tweets, total followers, and total followed accounts). After balancing, the estimates are computed using G-computation (66, 67) with robust standard errors. For comparability, estimates are standardized with effects interpreted in standard deviation changes.
@@ 4 Discussion
The American public is largely disinterested in politics (68) and so the aggregate consumption of news and political information is limited offline (31, 32, 33, 18), online (69) and on social media platforms (13, 70, 34, 41). And yet, because news consumption has many beneficial effects—ranging from increased political knowledge and participation, to more stable political attitudes, greater political tolerance, and higher support for democratic norms (19, 25, 71)—the question as to how to incentivize social media users to consume more quality news is important. Because most citizens go online for entertainment (38), this project proposed to link users’ online habits and interests with contextual nudges that could encourage them to consume more news and public affairs information on Twitter.
@@
Second, the bots presented as females led to more consistent and stronger increases in news following and the liking of news media content. Despite the fact that politics is still seen as a male domain (51, 52) and despite the mounting evidence that women are more likely than men to experience harassment or hate speech on social media (53, 54, 55), our sample was slightly more responsive to the intervention when it came from a female. Because Twitter users are predominantly male (61.2% vs 38.8%) (73) and because our sample was likely to have even more male users (given that sports was the dominant category), it is possible that men are more open to female bots nudging them and could feel potentially threatened by male bots telling them to follow news. Again, because these effects were small and not very robust, and because this evidence goes against the prevalent finding that politics and the online sphere are spaces where female opinion is disregarded, we encourage more work exploring these potentially differential reactions to pro-social online interventions coming from females versus males.
Third, the small detected effects were largely confined to the group of users who were already interested in politics. It is those who had previously tweeted about politics who showed the most substantial increases in the liking of news media contents as a result of our intervention (among those who were treated). In contrast, the bots failed to encourage users who were not interested in politics, as determined by their previous posting of political content on Twitter, to follow more news accounts, (re)tweet politics, or like news and political content.This finding speaks to the reinforcement hypothesis (29, 30) and suggests that people who are already politically interested are the ones who become yet more politically interested as a result of various interventions. Although incidental exposure to news and public affairs on social media platforms—which the comments from our bots effectively created—could serve an equalizing function and pull the previously disinterested citizens back into news and politics (44)—our study adds to the more pessimistic evidence that individuals with high political interest are becoming information richer and more participatory, whereas the ones with low political interest remain politically disengaged (29). We also note that whatever significant effects mostly emerged among the users who were tweeting about sports, potentially due to the fact that sports users were the majority of our sample. We caution against putting too much leverage on these findings because the sample sizes became substantially smaller when sub-divided into the three topic categories.
@@
In addition, it is not certain whether similar effects would emerge on a different platform, in a different time, or in a different sociopolitical context. Twitter is known to be an important channel for political information (74, 75), a key platform for politicians, journalists, and pundits (76), and one where many users express their political opinions and engage in political activities (77, 78). As such, engaging users with news and politics may have been more “natural” on and better integrated with Twitter than with Facebook or Instagram, where many users have more closely knit networks of friends and family. We encourage scholars to replicate our results on other platforms. Given the changes to the Twitter API after Elon Musk acquired Twitter, field experiments such as ours and others (79, 80, 28) might be no longer possible.
Lastly, our core focus was on encouraging the users to follow news accounts and engage with news and political information on social media, and so we cannot ascertain whether the treatment or the slightly increased news following and liking of news content had any effects on users’ political attitudes. A growing body of evidence suggest that although various (algorithmic) interventions can powerfully alter users’ on-platform exposures and behaviors, this has no corresponding effects on affective polarization, misperceptions, policy positions, among other attitudes (41, 81, 82, 83, 72). As aforementioned, however, our treatment could have triggered or enhanced political interest among some users, served as a gateway to hard news, or made users feel politically efficacious, outcomes that we did not measure and that are often overlooked in similar work.
@@
The API call we used to get the followed accounts was API.get_friends_ids(). The call for likes was API.get_favorites() and the call for (re-)tweets was API.user_timeline(). The final counts collected were as follows: followed accounts (pre N = 6,536,692, post N = 17,286,211), (re-)tweets (pre N = 2,285,401, post N = 2,201,009), and likes (pre N = 2,927,951, post N 2,846,354). More details can be found in SM. Incentivizing News Consumption on Social Media Platforms Using Large Language Models and Realistic Bot Accounts
@@
The authors gratefully acknowledge the support of the European Research Council, “Europeans exposed to dissimilar views in the media: investigating backfire effects,” Proposal EXPO- 756301 (ERC Starting Grant, Magdalena Wojcieszak – PI). The authors are also grateful to Zubair Shafiq and Muhammad Haroon for their research assistance. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the European Research Council.
[…]